\begin{thebibliography}{10}

\bibitem{BottTu}
R.~Bott and L.~W. Tu.
\newblock {\em Differential forms in algebraic topology}, volume~82 of {\em
  Graduate Texts in Mathematics}.
\newblock Springer-Verlag, New York-Berlin, 1982.

\bibitem{Hajlasz00}
P.~Haj\l{}asz.
\newblock Sobolev mappings, co-area formula and related topics.
\newblock In {\em Proceedings on {A}nalysis and {G}eometry ({R}ussian)
  ({N}ovosibirsk {A}kademgorodok, 1999)}, pages 227--254. Izdat. Ross. Akad.
  Nauk Sib. Otd. Inst. Mat., Novosibirsk, 2000.

\bibitem{HardtRiviere08}
R.~Hardt and T.~Rivi\`ere.
\newblock Connecting rational homotopy type singularities.
\newblock {\em Acta Math.}, 200(1):15--83, 2008.

\bibitem{Hatcher}
A.~Hatcher.
\newblock {\em Algebraic topology}.
\newblock Cambridge University Press, Cambridge, 2002.

\bibitem{Sucks}
K.~Mazowiecka and A.~Schikorra.
\newblock Minimal {$W^{s,\frac ns}$}-harmonic maps in homotopy classes.
\newblock {\em J. Lond. Math. Soc. (2)}, 108(2):742--836, 2023.

\bibitem{MS-stability}
K.~Mazowiecka and A.~Schikorra.
\newblock {$s$-stability for $W^{s, n/s}$-harmonic maps in homotopy groups}.
\newblock {\em Annales de l’Institut Henri Poincaré C, Analyse non linéaire
  (2024)}, published online first.

\bibitem{MonteilVanSchaftigen}
A.~Monteil and J.~Van~Schaftingen.
\newblock Uniform boundedness principles for sobolev maps into manifolds.
\newblock {\em Annales de l’Institut Henri Poincaré C, Analyse non
  linéaire}, 36(2):417–449, apr 2019.

\bibitem{ParkSchikorra}
W.~Park and A.~Schikorra.
\newblock Quantitative estimates for fractional {S}obolev mappings in rational
  homotopy groups.
\newblock {\em Nonlinear Anal.}, 235:Paper No. 113349, 16, 2023.

\bibitem{Riviere98}
T.~Rivi\`ere.
\newblock Minimizing fibrations and {$p$}-harmonic maps in homotopy classes
  from {$S^3$} into {$S^2$}.
\newblock {\em Comm. Anal. Geom.}, 6(3):427--483, 1998.

\bibitem{SacksUhlenbeck}
J.~Sacks and K.~Uhlenbeck.
\newblock The existence of minimal immersions of {$2$}-spheres.
\newblock {\em Ann. of Math. (2)}, 113(1):1--24, 1981.

\bibitem{Schikorra-VanSchaftingen-Hopf}
A.~Schikorra and J.~Van~Schaftingen.
\newblock An estimate of the {H}opf degree of fractional {S}obolev mappings.
\newblock {\em Proc. Amer. Math. Soc.}, 148(7):2877--2891, 2020.

\bibitem{VanSchaftingen2020}
J.~Van~Schaftingen.
\newblock {Estimates by gap potentials of free homotopy decompositions of
  critical Sobolev maps}.
\newblock {\em Advances in Nonlinear Analysis}, 9(1):1214--1250, 2020.

\bibitem{Whitehead_1947}
J.~H.~C. Whitehead.
\newblock {An Expression of Hopf's Invariant as an Integral}.
\newblock {\em Proceedings of the National Academy of Sciences of the United
  States of America}, 33(5):117--123, 1947.

\end{thebibliography}

\documentclass[12pt]{amsart}

\usepackage{tensor}     

\usepackage{tikz}
\usetikzlibrary{calc, decorations.pathreplacing, arrows.meta, patterns}

\usepackage[colorlinks=true,urlcolor=blue, citecolor=red,linkcolor=blue,linktocpage,pdfpagelabels, bookmarksnumbered,bookmarksopen]{hyperref}
\usepackage[hyperpageref]{backref}
\usepackage{cleveref}
\usepackage{xcolor}
\usepackage{amsthm} 
\usepackage{latexsym,amsmath,amssymb}
\usepackage{accents}
\usepackage[colorinlistoftodos,prependcaption,textsize=tiny]{todonotes}
\usepackage{a4wide}
\usepackage{soul}
\usepackage{mathtools} 
\usepackage{xparse} 
\usepackage{enumitem}		

\usepackage{calc}

\usepackage{accents}

\definecolor{indigo}{rgb}{0.29, 0.0, 0.51}
\definecolor{p1}{gray}{0.4}
\definecolor{p2}{gray}{0.6}
\definecolor{p3}{gray}{0.98}
\definecolor{p4}{gray}{0.8}
\definecolor{p5}{gray}{0.9}

plus 6pt minus 12pt
\def\eps{\varepsilon}

\def\B{{B}}

\def\N{{\mathbb N}}

\def\S{{\mathbb S}}

\newcommand{\degh}{\deg_{H}}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{definition}[theorem]{Definition}

\newcommand{\dif}{\,\mathrm{d}}
\newcommand{\dx}{\dif x}
\newcommand{\dy}{\dif y}
\newcommand{\dz}{\dif z}

\newcommand{\R}{\mathbb{R}}

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\C}{\mathbb{C}}

\newcommand{\brac}[1]{\left (#1 \right )}
\newcommand{\abs}[1]{\left |#1 \right |}

\newcommand{\la}{\mathopen{}\mathclose\bgroup\left\langle}
\newcommand{\ra}{\aftergroup\egroup\right\rangle}

\newcommand{\barint}{
\rule[.036in]{.12in}{.009in}\kern-.16in \displaystyle\int }

\newcommand{\barcal}{\mbox{$ \rule[.036in]{.11in}{.007in}\kern-.128in\int $}}

\def\mvint_#1{\mathchoice
          {\mathop{\vrule width 6pt height 3 pt depth -2.5pt
                  \kern -8pt \intop}\nolimits_{\kern -3pt #1}}%
          {\mathop{\vrule width 5pt height 3 pt depth -2.6pt
                  \kern -6pt \intop}\nolimits_{#1}}%
          {\mathop{\vrule width 5pt height 3 pt depth -2.6pt
                  \kern -6pt \intop}\nolimits_{#1}}%
          {\mathop{\vrule width 5pt height 3 pt depth -2.6pt
                  \kern -6pt \intop}\nolimits_{#1}}}

\numberwithin{theorem}{section} \numberwithin{equation}{section}

\newcommand{\aleq}{\precsim}

\def\XXint#1#2#3{{\setbox0=\hbox{$#1{#2#3}{\int}$}
     \vcenter{\hbox{$#2#3$}}\kern-.5\wd0}}

\let\latexchi\chi
\makeatletter
\renewcommand\chi{\@ifnextchar_\sub@chi\latexchi}
\newcommand{\sub@chi}[2]{
  \@ifnextchar^{\subsup@chi{#2}}{\latexchi^{}_{#2}}%
}
\newcommand{\subsup@chi}[3]{
  \latexchi_{#1}^{#3}%
}
\makeatother

\usepackage{mathtools} 

\title[Fractional harmonic maps from $\S^3$ to $\S^2$]{Existence of infinitely many homotopy classes\\ from $\S^3$ to $\S^2$ having a minimimzing $W^{s,\frac 3s}$-harmonic map}

\author{Adam Grzela}
\address[Adam Grzela]{Institute of Mathematics, %
University of Warsaw,
Banacha 2,
02-097 Warszawa, Poland}
\email{a.grzela@uw.edu.pl}

\author{Katarzyna Mazowiecka}
\address[Katarzyna Mazowiecka]{
Institute of Mathematics, %
University of Warsaw,
Banacha 2,
02-097 Warszawa, Poland}
\email{k.mazowiecka@mimuw.edu.pl}

\begin{document}
\subjclass[2010]{58E20, 35B65, 35J60, 35S05}
\begin{abstract}
  In 1998 T. Rivi\`{e}re proved that there exist infinitely many homotopy classes of $\pi_3(\S^2)$ having a minimizing 3-harmonic map. This result is especially surprising taking into account that in $\pi_3(\S^3)$ there are only three homotopy classes (corresponding to the degrees $\{-1,0,1\}$) in which a minimizer exists.

We extend this theorem in the framework of fractional harmonic maps and prove that for $s\in(0,1)$ there exist infinitely many homotopy classes of $\pi_{3}(\S^{2})$ in which there is a minimizing $W^{s,\frac{3}{s}}$-harmonic map.
 \end{abstract}
\maketitle
\sloppy

\maketitle
\tableofcontents
\sloppy

\section{Introduction}\label{s:introduction}
The homotopy groups \(\pi_{4m-1}(\mathbb{S}^{2m})\) for \(m \geq 1\) offer a rich source of computable invariants in homotopy theory. A classical construction due to Whitehead~\cite{Whitehead_1947} associates to each smooth map \(f \colon \mathbb{S}^{4m-1} \to \mathbb{S}^{2m}\) a numerical invariant known as the \emph{Hopf degree} (or \emph{Hopf invariant}). To define it, consider the standard volume form \(\omega_{\mathbb{S}^{2m}}\) on \(\mathbb{S}^{2m}\). The pullback \(f^* \omega_{\mathbb{S}^{2m}}\) is a closed \(2m\)-form on \(\mathbb{S}^{4m-1}\) and due to the Poincaré lemma this form is exact. Thus, there exists a \((2m-1)\)-form \(\eta \in \Omega^{2m-1}(\mathbb{S}^{4m-1})\) such that \(d\eta = f^* \omega_{\mathbb{S}^{2m}}\). The Hopf degree of \(f\) is then given by
\[
\degh(f) = \frac{1}{\left| \mathbb{S}^{2m} \right|^{2}} \int_{\mathbb{S}^{4m-1}} \eta \wedge d\eta.
\]
This definition is independent of the choice of \(\eta\), and the resulting quantity \(\degh\) is invariant under homotopy. For further properties of the Hopf invariant, we refer to~\cite{BottTu}.

In the case $m=1$ we have $\pi_{3}(\S^{2})\simeq \Z$ and $\degh(f)$ coincides with the homotopy class of $f$.

The Hopf invariant admits an estimate in terms of the critical Sobolev semi-norm; see~\cite[Introduction]{Riviere98}:
\begin{equation}\label{eq:hopf-estimate}
\left| \degh(f) \right| \lesssim \left( \int_{\mathbb{S}^{4m-1}} |\nabla f|^{4m-1} \right)^{\frac{4m}{4m-1}}.
\end{equation}
It is important to note that the exponent \(\frac{4m}{4m-1}>1\), and one might naturally ask whether this exponent could be improved. However, in the case \(m = 1\), Rivi\`{e}re showed in~\cite{Riviere98} that the exponent \(\frac{4}{3}\) in~\eqref{eq:hopf-estimate} is optimal. More precisely, he proved that
\begin{equation}\label{eq:optimalexponent}
\frac{\log \#_3 |d|}{\log d} \to \frac{3}{4} \quad \text{as} \quad |k| \to \infty,
\end{equation}
where for
\begin{equation}
 E_{1,3}(u)\coloneqq\int_{\mathbb{S}^3} |\nabla u|^3
\end{equation}
we define
\begin{equation}
\#_3 d \coloneqq \inf \left\{ E_{1,3}(u)\colon u \in W^{1,3}(\mathbb{S}^3, \mathbb{S}^2),\, \degh(u) = d \right\}.
\end{equation}
The optimality of an exponent greater than 1, together with the energy identity allowed Rivière to deduced:
\begin{theorem}[{\cite[Theorem I.1]{Riviere98}}]\label{th:Riviere}
There exist infinitely many homotopy classes of $\pi_{3}(\S^2)$ having minimizing $3$-harmonic maps, i.e., a map minimizing the $E_{1,3}$ energy in its homotopy class.
\end{theorem}
For completeness we state the energy identity in the spirit of \cite{SacksUhlenbeck}:
\begin{theorem}[Energy identity, {\cite[Proposition III.1]{Riviere98}}]\label{thm:identitylocal}
For all $d\in\Z$, there exists a finite sequence $(d_{1},\ldots,d_{N})$ of integers such that $\sum_{i=1}^N d_i = d$, and a finite collection of maps $v_{1},\ldots,v_{N}\colon\S^{3}\to \S^{2}$ of Hopf degrees $\degh(v_{i}) = d_{i}$, such that:
\begin{equation}\label{eq:localenergyidentity}
E_{1,3}(v_{i}) = \#_3d_{i} \quad \text{and} \quad \#_3 d = \sum_{i=1}^{N} E_{1,3}(v_{i}).
\end{equation}
\end{theorem}
 In order to prove \Cref{th:Riviere}, having \eqref{eq:optimalexponent} and \Cref{thm:identitylocal}, one can argue by contradiction. Suppose only finitely many homotopy classes admit minimizers.
 Then, letting $d_1$ denote the largest absolute value among these classes and $d_0$ the class with the smallest energy, the energy identity \eqref{eq:localenergyidentity} implies:
\begin{equation}
 \forall d\in\Z \quad \# d \ge \frac{\# d_0}{|d_1|} |d|,
\end{equation}
which contradicts \eqref{eq:optimalexponent} as $|d|\to\infty$.

In general, establishing optimality of exponents in estimates like \eqref{eq:hopf-estimate} appears to be a delicate question; see~\cite[Section 2.5.2 and Proposition 2.15]{HardtRiviere08} for related discussions. We note also that in the contradiction argument the precise value of the exponent did not play a role, we only used that it was not equal to 1.

Let now $s\in (0,1)$ and $p>1$. We denote by $W^{s,p}(\S^n)$ the Sobolev--Slobodeckij space with the semi-norm
\begin{equation}\label{eq:seminorm}
 E_{s,p}(u,\S^{n})= [u]_{W^{s,p}(\S^n)}^p\coloneqq \int_{\S^n}\int_{\S^n} \frac{|u(x) - u(y)|^p}{|x-y|^{n+sp}} \dx \dy.
\end{equation}
For $s\in (0,1)$, we define
\begin{equation}
\#_{s} \alpha \coloneqq \inf \left\{ E_{s,\frac ms}(u,\S^{n})\colon u \in W^{s,p}(\mathbb{S}^{n}, \mathbb{S}^{\ell}),\, \alpha \in \pi_{n}(\S^\ell) \right\}.
\end{equation}
In the case of elements of $\pi_3(\S^2)$ we write
\begin{equation}\label{eq:infimumoffractionalenergyhopf}
 \#_s d \coloneqq \inf \left\{ E_{s,\frac 3s}(u,\S^{3})\colon u \in W^{s,\frac 3s}(\mathbb{S}^{3}, \mathbb{S}^{2}),\, \deg_H(u)=d \right\}.
\end{equation}

In the fractional setting a counterpart to \eqref{eq:hopf-estimate} is possible and was established in \cite[Theorem 1.1]{Schikorra-VanSchaftingen-Hopf} for maps in $\pi_{4m-1}(\S^{2m})$ for the range $s\in[1-\frac{1}{4m}, 1)$
\begin{equation}\label{eq:fractioanlhopfestimate}
\abs{\deg_H(f)} \lesssim [f]^{\frac{4m}{s}}_{W^{s,\frac{4m-1}{s}}(\S^{4m-1})}.
\end{equation}
See also~\cite{ParkSchikorra} for related results. It remains an open problem to establish whether or not the fractional exponent range can be extended to $s\in(0,1)$.

Using Sobolev embedding, one can extend \Cref{th:Riviere} to fractional settings with $m=1$ and $s>\frac{3}{4}$. Indeed, by~\cite[Theorem 3.2]{MS-stability}, we have the following energy identity in the spirit of \cite{Sucks}:
\begin{theorem}[{\cite[Theorem 3.2]{MS-stability}}]\label{th:energy-id-frac}
Let $s\in(0,1)$, $n,\ell \in\N$, and assume that either $(n,\ell)=(1,1)$ or $\ell\ge 2$. For each $\alpha \in \pi_n\left(\mathbb{S}^{\ell}\right) \backslash\{0\}$ there exists a finite sequence $\left(\alpha_i\right)_{i=1}^N \subset \pi_n\left(\mathbb{S}^{\ell}\right) \backslash\{0\}$ such that
\begin{itemize}
\item[(1)] $\alpha=\sum_{i=1}^N \alpha_i$,
\item[(2)] $\#_{s} \alpha=\sum_{i=1}^N \#_{s} \alpha_i$,
\item[(3)] $\#_{s} \alpha_i$ are attained for each $i \in\{1, \ldots, N\}$.
\end{itemize}
\end{theorem}
Furthermore, the Sobolev embedding yields
\[
 [u]_{W^{s,\frac 3s}(\S^3)} \aleq \brac{\int_{\S^3} |\nabla u|^3}^\frac13
\]
which combined with \eqref{eq:optimalexponent} for large enough in absolute value $d$ gives the reverse inequality
\begin{equation}\label{eq:weakreverseinequality}
\#_{s} d \aleq \brac{\#_{3} d }^{\frac 1s}\aleq |d|^{\frac{3}{4s}}.
\end{equation}
For $s>\frac{3}{4}$ the exponent \eqref{eq:weakreverseinequality} is less than 1, allowing the same contradiction argument as in \Cref{th:Riviere}\footnote{Note that this is the same restriction on $s$ as in \cite{Schikorra-VanSchaftingen-Hopf}}.

The main result of this paper extends this existence result to the full range $s\in(0,1)$.
\begin{theorem}\label{th:main}
Let $s\in(0,1)$. There exist infinitely many homotopy classes of $\pi_{3}(\S^{2})$ having a minimizing $W^{s,\frac{3}{s}}(\S^{3},\S^{2})$-harmonic map.
\end{theorem}
In particular, we obtain that for $m=1$ the exponent in \eqref{eq:fractioanlhopfestimate} is optimal, see \Cref{3-3-thm:asymptote}.

\subsection*{Notation}

A map \( u \) is said to be a \emph{minimizing} \( W^{s,\frac{3}{s}}(\mathbb{S}^{3},\mathbb{S}^{2}) \)-\emph{harmonic map} if it minimizes the energy \( E_{s,\frac{3}{s}} \) within its homotopy class. For any open set \( \Omega \subset \mathbb{S}^{n} \), we define
\[
E_{s,p}(u,\Omega) \coloneqq \int_\Omega \int_\Omega \frac{|u(x) - u(y)|^p}{|x - y|^{n + sp}} \, \mathrm{d}x \, \mathrm{d}y.
\]

We denote by \( B(x,r) \) the geodesic ball in \( \mathbb{S}^m \) centered at \( x \) with radius \( r \). When the center is irrelevant, we simply write \( B(r) \). We use the notation \( A \lesssim B \) to indicate that there exists a constant \( C \) such that \( A \leq C B \), where \( C \) is independent of any essential parameter.

{\bf Acknowledgment.} The project is co-financed by
\begin{itemize}
 \item (AG, KM) the Polish National Agency for Academic Exchange within Polish Returns Programme -
BPN/PPO/2021/1/00019/U/00001;
\item (KM) the National Science Centre, Poland grant No. 2023/51/D/ST1/02907.
\end{itemize}

\section{Prerequisites}
To prepare for the proof of \Cref{th:main}, we begin by recalling the key properties of the classical Hopf map $h\colon \S^3 \to \S^2$, which will play a central role in our construction. We also collect several auxiliary tools that will allow us to localize energy, glue maps, and modify them homotopically while controlling the Sobolev energy.
\begin{definition} The Hopf map $h\colon \S^3\to\S^2$ is defined by:
\begin{equation}\label{eq:Hopfmap}
\begin{split}
 h\colon \S^3 \subset\R^4\simeq \C\times \C \to \S^2 \subset \R^3\simeq \R\times\C \\
 h(w,z) = (|w|^2 - |z|^2,2w\overline{z}).
 \end{split}
\end{equation}
\end{definition}
We will use the following properties:
\begin{enumerate}[label=(\Alph*)]
 \item $\deg_H(h) = 1$ (see, e.g., \cite[Example 4.45]{Hatcher});
 \item\label{item:hopfproperties} $|h^\ast \omega_{\S^2}|=\frac{1}{2}|\nabla h|^2 = 4$.
\end{enumerate}
The following technical lemmas will be used to estimate the fractional Sobolev energy of maps defined on unions of sets or modified locally. These are essential for adapting Rivière’s construction to the fractional framework.

The first lemma allows us to localize the $W^{s,p}$-energy of a map with a "buffer zone".
\begin{lemma}[{\cite[Lemma 2.2]{MonteilVanSchaftigen}}]\label{le:gluing}
Let $s\in (0,1]$, $1\leq p< \infty$, and let $\mathcal N$ be a connected Riemannian manifold. There exists a constant $C= C(s,p,m)>0$,  such that for every $\eta\in (0,1)$, every open set $A\subset \R^m$, every measurable function $u\colon A\to \mathcal N$, and every $\rho>0$ such that $\B(\rho)\setminus \overline{B(\eta\rho)}\subset A$,
\begin{equation}
E_{s,p}(u,A) \leq \left( 1 + \frac{C}{(1-\eta)^{sp+1}} \right)E_{s,p}(u,B(\rho)) +
\left( 1 + \frac{C\eta^{m}}{1-\eta} \right) E_{s,p}(u,A\setminus B(\eta\rho)).
\end{equation}
\end{lemma}

We use the next lemma for connecting together maps defined on a finite (or potentially countably infinite) number of small patches:
\begin{lemma}[{\cite[Lemma 2.3]{MonteilVanSchaftigen}}]\label{le:patching}
Let $s\in (0,1)$, $1\leq p< \infty$. Let $I$ be a finite or countably infinite set and for each $i\in I$ let $u_i\colon \S^3 \to \S^2 $ be a measurable map. If there exists $b\in \S^2 $ and a collection $(A_i)_{i\in I}$ of open subsets of $\S^3 $ such that $A_i\cap A_j = \emptyset$ for $i\neq j$ and $u_i(x)\equiv b$ for $x\in \S^3 \setminus A_i$, then for a map defined as:
\begin{equation}
u(x) =
\begin{cases}
u_i(x) &\text{ if } x\in A_i \\
b &\text{otherwise}
\end{cases}
\end{equation}
we have
\begin{equation}
E_{s,p}(u, \S^3 ) \leq 2^p \sum_{i\in I} E_{s,p}(u_i, \S^3 ).
\end{equation}
\end{lemma}

The last lemma will allow us to "open" maps, that is produce a homotopic map which is locally constant, equal to some chosen point. Its corollary states, that we retain control over the $W^{s,p}$-energy of such an "opened" map.
\begin{lemma}[{\cite[Lemma 3.4]{VanSchaftingen2020}}]\label{le:holeing}
For every $b \in \S^2 $ and every $\varepsilon>0$, there exists a map $\Theta \in C^1(\S^2 , \S^2 )$ which is homotopic to the identity and such that $\Theta\equiv b$ in a neighborhood of $b$ and for every $x, y \in \S^2 $, we have
$d_{\S^2 }(\Theta(x), \Theta(y))\leq (1+\varepsilon)d_{\S^2 }(x,y)$.
\end{lemma}

\begin{corollary}\label{cor:energy-holeing}
In particular, for any given $u\colon \S^3\to\S^2$ and $b\in u(\S^3)\subset \S^2$, any $\eps>0$, we have for $\Theta$ from \Cref{le:holeing} 
\begin{equation}
E_{s,p}(\Theta\circ u, \S^3 ) \leq (1+\varepsilon)^pE_{s,p}(u,\S^3 ),
\end{equation}
$\Theta\circ u \sim u$ and $\Theta\circ u$ is locally constant (on the neighborhood of the fiber $u^{-1}(b)$).
\end{corollary}

\section{Proof of \Cref{th:main}}
As explained in the introduction, the key to proving \Cref{th:main} lies in showing that the exponent in estimate (1.9) is optimal in the case $m=1$. This is the content of the following theorem:
\begin{theorem}\label{3-3-thm:asymptote}
Let $sp = 3$. There exists a constant $C=C(s)>0$ such that
\begin{equation}
\#_{s} d \leq C |d|^{\frac{3}{4}}
\end{equation}
where $\#_{s} d$, defined in \eqref{eq:infimumoffractionalenergyhopf}, denotes the infimum of $E_{s,\frac3s}$ among the maps
$u\colon\S^3 \to \S^2$ of Hopf degree $d$.
\end{theorem}

\begin{proof}
We follow the construction from the proof of \cite[Lemma III.1]{Riviere98}, and show that it also works in the nonlocal setting for fractional harmonic maps. For readability we introduce $p=\frac 3s$.

\textsc{Step 1. Proof in the case $d=k^2$}

 Consider the Hopf fibration $h\colon\S^3 \to\S^2$ and a map $v\colon\S^2\to\S^2$ of topological degree $\deg v = k$. Recall that (see, e.g., \cite[p.436]{Riviere98}):
\begin{equation}
\degh(v\circ h) = (\deg v)^2 \degh h = k^2.
\end{equation}
By the co-area formula, see, e.g., \cite{Hajlasz00}, we have:
\begin{equation}
\begin{split}
E_{s,p}(v\circ h, \S^3 ) &= \int_{\S^3 } \int_{\S^3 } \frac{|v\circ h(x) - v\circ h(y)|^p}{|x-y|^{3+sp}} \dx\dy \\
&= \int_{\S^3 } \int_{\S^3 } \frac{|v\circ h(x) - v\circ h(y)|^p}{|x-y|^{3+sp}} \frac{|h^*\omega_{\S^2}|}{|h^*\omega_{\S^2}|}\frac{|h^*\omega_{\S^2}|}{|h^*\omega_{\S^2}|}\dx\dy\\
&= \frac{1}{16}\int_{\S^2} \int_{\S^2} \left(\int_{h^{-1}(z_2)}\int_{h^{-1}(z_1)}\frac{|v\circ h(x) - v\circ h(y)|^p}{|x-y|^{3+sp}}\dx\dy \right) \dz_1\dz_2,
\end{split}
\end{equation}
in the last equality we used \ref{item:hopfproperties}. Now, since $h$ is constant on its fibers, we have:
\begin{equation}\label{eq:1}
\begin{split}
E_{s,p}(v\circ h, \S^3 ) &=
\frac{1}{16}\int_{\S^2} \int_{\S^2} \left(\int_{h^{-1}(z_2)}\int_{h^{-1}(z_1)}\frac{|v\circ h(x) - v\circ h(y)|^p}{|x-y|^{3+sp}}\dx\dy \right) \dz_1\dz_2 \\
&=\frac{1}{16}\int_{\S^2} \int_{\S^2} |v(z_1) - v(z_2)|^p \left(\int_{h^{-1}(z_2)}\int_{h^{-1}(z_1)}\frac{1}{|x-y|^{3+sp}}\dx\dy \right) \dz_1\dz_2.
\end{split}
\end{equation}
In order to reduce the exponent by one we integrate the kernel $|x-y|^{-3-sp}$ along the 1-dimensional fibers. Rewrite the integral using the Cavalieri principle:
\begin{equation}\label{eq:2}
\begin{split}
\int_{h^{-1}(z_2)}\int_{h^{-1}(z_1)}\frac{1}{|x-y|^{3+sp}}\dx\dy =
\int_{h^{-1}(z_2)}\left(\int_{S}\frac{1}{|t|^{3+sp}}\dif\mathcal{H}^1(t) \right) \dy  \\
= \int_{h^{-1}(z_2)}\left(\int_{0}^{\infty}(3+sp) \tau^{2+sp} \mathcal{H}^1\left(S \cap \left\{t:\ \frac{1}{|t|}> \tau\right\}\right) \dif\tau \right) \dy ,
\end{split}
\end{equation}
where $S = -y + h^{-1}(z_1)$ is a circle of radius 1 centered around the point $-y$. Now, notice that:
\begin{equation}\label{3-1-eq:2pitau}
\mathcal{H}^{1}\left(S \cap \left\{t:\ \frac{1}{|t|}> \tau\right\}\right) =
\mathcal{H}^{1}\left(S \cap \left\{t:\ |t|< \frac{1}{\tau}\right\}\right) \leq \frac{2\pi}{\tau},
\end{equation}
which follows from the fact, that a ball cannot contain a circle arc longer than $2\pi$ times its radius. Additionally, observe, that since $|z_1 - z_2|\leq \text{Lip}(h) |x - y| = \text{Lip}(h)|t|$:
\begin{equation}\label{3-1-eq:bound}
\mathcal{H}^{1}\left(S \cap \left\{t:\ |t|< \frac{1}{\tau}\right\}\right) = 0 \text{ for } \tau \geq \frac{\text{Lip}(h)}{|z_1 - z_2|}.
\end{equation}
Putting (\ref{3-1-eq:2pitau}) and (\ref{3-1-eq:bound}) together, we obtain:
\begin{equation}\label{eq:3}
\begin{split}
\int_{0}^{\infty}(3+sp) \tau^{2+sp} \mathcal{H}^1\left(S \cap \left\{t:\ \frac{1}{|t|}> \tau\right\}\right) \dif\tau &\leq \int_0^{\frac{\text{Lip}(h)}{|z_1-z_2|}}(3+sp)\tau^{2+sp}\frac{2\pi}{\tau} \dif\tau \\
&\lesssim \frac{1}{|z_1 - z_2|^{2+sp}}.
\end{split}
\end{equation}
Consequently combining \eqref{eq:1} with \eqref{eq:2} and \eqref{eq:3} we obtain
\begin{equation}
E_{s,p}(v\circ h,\S^3)\aleq \int_{\S^2}\int_{\S^2} \frac{|v(z_1)-v(z_2)|^p}{|z_1 - z_2|^{2+sp}}\dif z_1 \dif z_2 = E_{s,p}(v,\S^2).
\end{equation}
Thus we conclude, that:
\begin{equation}\label{3-1-eq:poczatek}
\#_{s}k^2 \leq E_{s,p}(v\circ h, \S^3 ) \lesssim E_{s,p}(v,\S^2)
\end{equation}
for any $v\colon\S^2\to\S^2$ of degree $k$.

In particular, we may take \( v \) from the construction in~\cite[p.~437]{Riviere98}, which we briefly recall here. There exists a constant \( \lambda > 0 \) such that, for any \( k \in \mathbb{N} \), one can find \( k \) disjoint geodesic balls \( \left(B(x_i, \frac{\lambda}{\sqrt{k}})\right)_{i=1}^k \) of radius \( \frac{\lambda}{\sqrt{k}} \) in \( \mathbb{S}^2 \). For each \( i \), let \( v_i \colon \mathbb{S}^2 \to \mathbb{S}^2 \) be a degree-one map that is constant (equal to some fixed \( b \in \mathbb{S}^2 \)) outside of \( B(x_i, \frac{\lambda}{\sqrt{k}}) \), and satisfies \( |\nabla v_i| \leq C\sqrt{k} \). For a detailed construction, we refer the reader to \Cref{a1-3-lmm:top-deg1}. We then define \( v \colon \mathbb{S}^2 \to \mathbb{S}^2 \) by
\begin{equation}\label{eq:map-construction}
v = \begin{cases}
v_i & \text{on } B(x_i, \frac{\lambda}{\sqrt{k}}), \\
b & \text{otherwise}.
\end{cases}
\end{equation}

By construction,
\begin{equation}\label{eq:map-properties}
\deg v = k \quad \text{and} \quad |\nabla v|\leq C\sqrt{k}
\end{equation}
since it is a local property.

Employing \Cref{le:patching}, we write:
\begin{equation}\label{eq:firstvestimate}
E_{s,p}(v, \S^2) \leq 2^p\sum_{i=1}^kE_{s,p}(v_i, \S^2).
\end{equation}
Additionally, by using Lemma \ref{le:gluing} for $\eta = 1/2$ and $\rho = \frac{2\lambda}{\sqrt{k}}$, we obtain for some $C=C(s)$ (in particular the constant is independent of $k$):
\begin{equation}
\begin{split}
E_{s,p}\left(v_i, \S^2\right) &\leq
C E_{s,p}\left(v_i, B(x_i,\frac{2\lambda}{\sqrt{k}})\right) +
C E_{s,p}\left(v_i,\S^2\setminus B(x_i,\frac{\lambda}{\sqrt{k}})\right) \\
&\lesssim E_{s,p}\left(v_i,  B(x_i,\frac{2\lambda}{\sqrt{k}})\right),
\end{split}
\end{equation}
where the last inequality is a consequence of $v_i$ being constant on the set $\S^2\setminus B(x_i,\frac{\lambda}{\sqrt{k}})$. It remains to estimate the localized energy:
\begin{equation}\label{3-1-eq:koniec}
\begin{split}
E_{s,p}\left (v_i,  B(x_i,\frac{2\lambda}{\sqrt{k}})\right) &=
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\frac{|v_i(x) - v_i(y)|^p}{|x-y|^{2+sp}}\dx\dy\\
&=
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\frac{|v_i(x) - v_i(y)|^p}{|x-y|^{p}} \frac{1}{|x-y|^{2+(s-1)p}} \dx\dy \\
&\lesssim (\sqrt{k})^{p}
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\frac{1}{|x-y|^{2+(s-1)p}}\dx\dy \\
&\leq k^{p/2}
\int_{ B(x_i,\frac{2\lambda}{\sqrt{k}})}
\int_{ B(0,\frac{4\lambda}{\sqrt{k}})}
\frac{1}{|z|^{2+(s-1)p}} \dz \dy\\
&\lesssim k^{p/2} \left( \frac{1}{\sqrt{k}} \right)^2 \left( \frac{1}{\sqrt{k}} \right)^{-(s-1)p} = k^{\frac{sp}{2} - 1}.
\end{split}
\end{equation}
Putting \eqref{3-1-eq:poczatek} with \eqref{eq:firstvestimate}-\eqref{3-1-eq:koniec} together with the fact, that $sp=3$, we obtain:
\begin{equation}
\#_{s}k^2\lesssim E_{s,p}(v, \S^2) \lesssim \sum_{i=1}^k E_{s,p}(v_i, \S^2) \lesssim k^{sp/2} = \left(k^2\right)^{3/4},
\end{equation}
which concludes the first part of the proof.

\textsc{Step 2: Proof for any $d\in\Z$}

To prove the claim for arbitrary Hopf degree, fix any \( d \in \mathbb{Z} \setminus \{0\} \). Without loss of generality, assume \( d > 0 \), as the case \( d < 0 \) follows by precomposing with an orientation-reversing diffeomorphism of \( \mathbb{S}^3 \), which changes the sign of the Hopf invariant. Choose \( k \in \mathbb{N} \) such that
\[
k^2 < d \leq (k+1)^2.
\]
Then
\[
0 < d - k^2 < 2k < 2\sqrt{d},
\]
which will be useful in the estimates below.

Given any map $w\colon \S^3 \to \S^2$ with $\degh(w)=k^2$, pick a point $b\in w(\S^3)\subset \S^2$, take an $\eps>0$, and use Lemma \ref{le:holeing} and Corollary \ref{cor:energy-holeing}, we obtain another map $\tilde{w} \sim w$, which is locally constant (equal to $b$) in the neighborhood of the fiber $w^{-1}(b)$ and satisfies relevant energy bounds, i.e.,
\begin{equation}\label{3-3-eq:control}
E_{s,p}(\tilde{w}, \S^3 ) \leq (1+\varepsilon)^pE_{s,p}(w,\S^3 ).
\end{equation}
As such, $\tilde{w}^{-1}(b)$ contains a non-empty open set and we can find $d-k^2$ disjoint open balls $\left( B(x_i,r)\right)_{i=1}^{d-k^2}$ of positive radius inside it.

We can use Hopf degree one maps $g_{x_i,r}$ with respective supports in $B(x_i,r)$ (i.e. constant equal to $b$ outside a given ball), $E_{s,p}(g_{x_i,r},\S^3 ) = E$ for $i=1,\ldots,d-k^2$ (see: Corollary \ref{a1-4-cor:hopf-deg1}), to define $u\colon \S^3\to\S^2$  by:
\begin{equation}
u(x) = \begin{cases}
g_{x_i,r}(x) &\text{if } x\in  B(x_i,r)\\
\tilde{w}(x) &\text{otherwise}.
\end{cases}
\end{equation}
Note, that $\degh(u)=d$ by construction. Additionally, observe that $u$ satisfies the assumptions of \Cref{le:patching}. We have a finite collection of nonempty, nonintersecting open sets $(A_i)_{i=0}^{d-k^2}$, namely:
\begin{equation}
A_0 = \S^3 \setminus \tilde{w}^{-1}(b), \quad A_i =  B(x_i,r),
\end{equation}
and a collection of measurable maps, all constant equal to a common point $b$ outside their specified domain: $\tilde{w}$ for $A_0$ and $g_{x_i,r}$ for each of $A_i$.

Therefore, by Lemma \ref{le:patching}, we have:
\begin{equation}
\#_{s}d\leq E_{s,p}(u,\S^3 )\lesssim
E_{s,p}(\tilde{w}, \S^3 ) + (d-k^2)E
\lesssim (1+\eps)^p E_{s,p}(w, \S^3 ) + (d-k^2).
\end{equation}
Where in the last inequality we used \eqref{3-3-eq:control}. Passing to the infimum on the right-hand side and using the arbitrariness of $\eps>0$, we obtain:
\begin{equation}
\#_{s}d \lesssim \#_{s}k^2 + (d-k^2) \lesssim d^{3/4} + d^{1/2} \lesssim d^{3/4},
\end{equation}
which concludes the proof.
\end{proof}

We now combine the optimal upper bound from \Cref{3-3-thm:asymptote} with the energy decomposition of \Cref{th:energy-id-frac} to conclude the proof of our main result.
\begin{proof}[Proof of Theorem \ref{th:main}]
Suppose only finitely many homotopy classes of $\pi_3(\S^2 )$ contain a minimizing $W^{s,\frac{3}{s}}$-harmonic map. Let $\{d_i\}_{i=0}^K$ be the integers corresponding via the Hopf degree to these homotopy classes, with infima of energies again denoted as $\#_{s}{d_i}$. Now, we may suppose (up to reordering), that $\#_{s}{d_{0}}$ is the smallest of those energies and $|d_{1}|$ is the largest of the integers. By \Cref{th:energy-id-frac} we get for any $d$ a sequence $(a_i)\subset \{1,\ldots,K\}$ such that $d = \sum d_{a_i}$ and:
\begin{equation}
\#_{s}{d} = \sum_{i=1}^{N} \#_{s}{d_{a_i}} =
\sum_{i=1}^{N} \frac{|d_{a_i}|}{|d_{a_i}|}\#_{s}{d_{a_i}} \geq
\#_{s}{d_{0}}\sum_{i=1}^{N} \frac{|d_{a_i}|}{|d_{a_i}|} \geq
\frac{\#_{s}{d_{0}}}{|d_{1}|}\sum_{i=1}^{N} |d_{a_i}| \geq \frac{\#_{s}d_0}{|d_1|}|d|
\end{equation}
which, as $|d|$ tends to infinity, clearly contradicts Theorem \ref{3-3-thm:asymptote}.
\end{proof}

\appendix
\section{}
\begin{lemma}\label{a1-3-lmm:top-deg1}
Let \( x_0, b \in \mathbb{S}^m \) and \( 0 < r < 1 \). There exists a differentiable map \( f_{x_0,r} \colon \mathbb{S}^m \to \mathbb{S}^m \) such that
\[
\deg(f_{x_0,r}) = 1, \qquad |\nabla f_{x_0,r}| \leq \frac{C}{r}, \qquad f_{x_0,r} \equiv b \text{ outside } B(x_0, r),
\]
for some constant \( C > 0 \) independent of \( r \).
\end{lemma}

\begin{proof}
Let \( x_N = (0,\ldots,1) \) and \( x_S = (0,\ldots,-1) \) be the north and south poles of \( \mathbb{S}^m \), and let \( \Pi \colon \mathbb{S}^m \setminus \{x_N\} \to \mathbb{R}^m \) denote the stereographic projection, with inverse
\[
\Pi^{-1}(y) = \left( \frac{2y}{1 + |y|^2}, \frac{-1 + |y|^2}{1 + |y|^2} \right).
\]
By rotating the domain and codomain, we may assume \( x_0 = x_S \) and \( b = x_N \).

\textit{Step 1.} Define the dilation \( \mu_r(y) = y/r \), and consider the map
\[
\tilde{f} = \Pi^{-1} \circ \mu_r \circ \Pi.
\]
Then \( \tilde{f} \) maps \( B(x_S, r) \) onto the southern hemisphere and satisfies \( |\nabla \tilde{f}| \leq C/r \) on \( B(x_S, r) \), since \( \Pi \) and \( \Pi^{-1} \) are smooth diffeomorphisms on this region.

\textit{Step 2.} Let \( g \colon \mathbb{S}^m \to \mathbb{S}^m \) be the smooth map
\[
g(x) = (-2x_1x_{m+1}, \ldots, -2x_mx_{m+1}, 1 - 2x_{m+1}^2),
\]
which maps the southern hemisphere onto \( \mathbb{S}^m \) and satisfies \( g(x) = x_N \) on the equator. One verifies \( |\nabla g| \leq C \).

Define
\[
f_{x_S, r}(x) =
\begin{cases}
g(\tilde{f}(x)) & \text{if } x \in B(x_S, r), \\
x_N & \text{otherwise}.
\end{cases}
\]
Then \( f_{x_S, r} \) has degree 1 and satisfies the desired bounds. Composing with rotations yields the general case \( f_{x_0,r} \), completing the proof.
\end{proof}

\begin{corollary}\label{a1-4-cor:hopf-deg1}
Let \( x_0 \in \mathbb{S}^3 \), \( b \in \mathbb{S}^2 \), and suppose \( sp = 3 \), \( 0 < r < 1 \). Then there exists a differentiable map \( g_{x_0, r} \colon \mathbb{S}^3 \to \mathbb{S}^2 \) such that
\[
\degh(g_{x_0,r}) = 1, \qquad E_{s,p}(g_{x_0,r}, \mathbb{S}^3) \leq E, \qquad g_{x_0,r} \equiv b \text{ outside } B(x_0, r),
\]
for some constant \( E > 0 \) independent of \( r \).
\end{corollary}

\begin{proof}
Let \( h \colon \mathbb{S}^3 \to \mathbb{S}^2 \) be the Hopf map, and choose any \( b' \in h^{-1}(b) \). Let \( f_{x_0,r} \colon \mathbb{S}^3 \to \mathbb{S}^3 \) be the map from \Cref{a1-3-lmm:top-deg1}, which equals \( b' \) outside \( B(x_0, r) \), and define
\[
g_{x_0,r} = h \circ f_{x_0,r}.
\]
Then \( g_{x_0,r} \equiv b \) outside \( B(x_0, r) \), and
\[
\degh(g_{x_0,r}) = \degh(h) \cdot \deg(f_{x_0,r}) = 1.
\]

Since \( g_{x_0,r} \) is constant outside \( B(x_0, r) \), Lemma~\ref{le:gluing} with \( \eta = 1/2 \), \( \rho = 2r \) gives
\[
E_{s,p}(g_{x_0,r}, \mathbb{S}^3) \leq C\, E_{s,p}(g_{x_0,r}, B(x_0,2r)).
\]
Using the Lipschitz continuity of \( h \) and the estimate \( |\nabla f_{x_0,r}| \leq C/r \), we compute
\begin{align*}
E_{s,p}(g_{x_0,r}, B(x_0,2r)) 
&\leq C \int_{B(x_0,2r)} \int_{B(x_0,2r)} 
\frac{|f_{x_0,r}(x) - f_{x_0,r}(y)|^p}{|x - y|^{3 + sp}} \,dx\,dy \\
&\leq \frac{C}{r^p} \int_{B(x_0,2r)} \int_{B(0,4r)} \frac{1}{|z|^{3 + (s-1)p}} \,dz\,dy \\
&\leq \frac{C}{r^p} \cdot r^3 \cdot r^{-(s-1)p} = C,
\end{align*}
where we used \( sp = 3 \) in the final step. This concludes the proof.
\end{proof}

\bibliographystyle{abbrv}%
\bibliography{bib}%
\end{document}